# Surrogate-based multilevel Monte Carlo methods for uncertainty quantification in the Grad-Shafranov free boundary problem

### v2

@@ 1.1 The Grad-Shafranov free boundary problem with uncertainty
In a magnetic confinement fusion reactor, strong magnetic fields produced by external coils are used to contain a hot plasma in the interior of a vacuum vessel. The confined plasma—a mixture of ionized light atomic nuclei—is heated through electromagnetic radiation to induce the merging of the atomic nuclei and the release of a significant amount of energy in a process known as a thermonuclear reaction. During confinement, charged particles in the plasma undergo forces due to the hydrostatic pressure, $p$, and to the magnetic field, $\boldsymbol{B}$, generated by the currents, $\boldsymbol{J}$, flowing through the external coils and through the plasma. A state of equilibrium is reached when the magnetic pressure equals the hydrostatic pressure and is expressed mathematically by requiring that the condition ${\boldsymbol{J}}\times{\boldsymbol{B}}=\nabla p$ holds in the entire space.
If the reactor possesses axial symmetry, the equilibrium condition can then be stated in terms of a nonlinear partial differential equation for a scalar variable known as the poloidal flux function, $u$. In this study, we will focus on a particular family of axially symmetric reactors known as tokamaks. The cross section of the typical configuration is depicted schematically in Figure 1. A vacuum vessel (solid curved line) contains the plasma (violet region), while an array of coils (solid gray rectangles) generate the confining magnetic field.
@@
$-\nabla\,\cdot\,\left(\frac{1}{\mu r}\nabla u\right)=J_{\varphi}(u,r),$
[…]
@@
    Within the plasma (confined within the region $\Omega_{p}$), the current can be represented in terms of the hydrostatic pressure and an additional scalar variable known as the toroidal field $g$ as
    $J_{\varphi}(u,r)=r\frac{d}{du}p(u)+\frac{1}{2\,\mu r}\frac{d}{du}g^{2}(u).$
Both the pressure and the toroidal field are considered constant along the streamlines of the magnetic field and thus take the forms $p=p(u)$ and $g=g(u)$. In this study, following the model proposed in [30], we adopt the following forms
$\frac{d}{du}p(u)=j_{0}\frac{\beta}{r_{0}}\left(1-u_{N}^{a_{1}}\right)^{a_{2}},\qquad\text{ and }\qquad\frac{1}{2}\frac{d}{du}g^{2}(u)=j_{0}\mu_{0}r_{0}(1-\beta)\left(1-u_{N}^{a_{1}}\right)^{a_{2}},$ (1)
where $u_{N}\in[0,1]$ represents a normalization of $u$. The parameter $r_{0}$ corresponds to the outer radius of the vacuum chamber, while $a_{1}$ and $a_{2}$ govern the sharpness of the current peaks near the magnetic axis; $\beta$ referred to as poloidal beta, serves as a parameter measuring the ratio between the hydrostatic pressure in the plasma due to temperature (usually called plasma pressure) and the pressure attributed to the external magnetic field (or magnetic pressure), $\mu_{0}$ is the magnetic permeability of vacuum, and $j_{0}$ is introduced as a normalization factor. The region $\Omega_{p}$ circumscribed by the last closed level set of $u$ that does not intersect any component of the reactor determines the boundary of the plasma, $\partial\Omega_{p}$, and thus the confinement region. When $\partial\Omega_{p}$ goes through a saddle point of $u$, it is referred to as a separatrix, and the saddle point is known as an x-point. Since $u$ and $\Omega_{p}$ are not known a priori this leads to a free boundary problem. Moreover, the dependence of the source terms $p(u)$ and $g(u)$—and possibly even of the magnetic permeability $\mu$—on the solution add further complexity to the Grad-Shafranov free boundary problem.
In practice, all the parameters of the model are susceptible to uncertainties (arising from imperfect measurements, operational variations, engineering tolerances, etc.) that introduce stochasticity into the solution $u$ and its derived quantities. In this study we will consider that the uncertainties may affect the parameters appearing in the definition of the source term $J_{\varphi}$; namely, the current intensities of the external coils, and the parameters appearing in (1). To model these uncertainties, we introduce a $d$-dimensional random variable $\boldsymbol{\omega}:=(\omega_{1},\ldots,\omega_{d})$, whose components are independent and uncorrelated, and whose $k$th component $\omega_{k}$ is associated with the uncertainty in the $k$th parameter. We will denote the joint probability density function by $\pi(\boldsymbol{\omega})$ and the parameter space by $\boldsymbol{W}$.
@@
$-\nabla\,\cdot\,\left(\frac{1}{\mu(u,\boldsymbol{\omega})r}\nabla u(\cdot,\boldsymbol{\omega})\right)=\left\{\begin{array}[]{ll}r\frac{d}{du}p(u,\boldsymbol{\omega})+\frac{1}{2\,\mu r}\frac{d}{du}g^{2}(u,\boldsymbol{\omega})&\quad\text{ in }\Omega_{p}(u,\boldsymbol{\omega})\,,\\ I_{k}(\boldsymbol{\omega})/S_{k}&\quad\text{ in }\Omega_{W_{k}}\\ 0&\quad\text{ elsewhere.}\end{array}\right.$ (2)
@@
[…]
@@ 1.2 Overview of the article
In its simplest form, a Monte Carlo estimate of the expectation $\mathbb{E}[u]$ gathers a large number, $N$, of the realizations of $u(\cdot,\boldsymbol{\omega})$ and estimates the expectation by the mean of the sample as
@@ 2.1 Sparse grid stochastic collocation
We now present a brief overview of the sparse grid stochastic collocation method [3, 27, 31, 36] for approximating the solution to (2) with stochastic parameters. Assume that for every fixed realization of the parameter $\boldsymbol{\omega}\in\boldsymbol{W}$, $u(\cdot,\boldsymbol{\omega})$ is an element of a function space $Z$. We wish to build an interpolatory approximation of $u(\cdot,\boldsymbol{\omega})$ based on a set of observations at particular values of $\boldsymbol{\omega}$. For simplicity, we suppose that the values of $x$ are restricted to the $d$-dimensional unit cube $[0,1]^{d}$. The stochastic collocation method uses particular values of $\boldsymbol{\omega}$ that give rise to a nested family of grids in the parameter space, where each grid consists of $m_{i}$ values of $\boldsymbol{\omega}$ that we shall call nodes. To mitigate the so-called curse of dimensionality, the sequence of grids can be built by choosing the nodes according to Smolyak’s rule, which curtails the growth in the number of nodes involved—although at the expense of some accuracy—and leads to a sequence of grids known as sparse grids. The construction process can be summarized as follows. Let
@@ 2.2 Notational considerations
In our context, $Z$ is the solution space to the free boundary problem (2) defined in [21] as
[…]
where $\Omega$ denotes a bounded, Lipschitz region, fully enclosing all the external coils $\Omega_{c_{i}}$ and the component structures of the reactor. The inner product and energy norm for this space are defined as
[…]
@@
[…]
@@ 2.3 Collocation with a single level of spatial discretization
$\mathcal{W}_{\text{SL-SL}}^{\text{off}}=PW\simeq\epsilon^{-\frac{1}{\nu}}\cdot\epsilon^{-\frac{\gamma}{\alpha}}=\epsilon^{-\frac{1}{\nu}-\frac{\gamma}{\alpha}},$ (14)
@@ 2.4 Collocation with multilevel spatial discretization
[…]
@@
[…]
@@
$\sum_{\ell=0}^{L}W_{\ell}\simeq\sum_{\ell=0}^{L}{M_{\ell}^{\gamma}}=M_{0}^{\gamma}\left(\sum_{\ell=0}^{L}s^{\gamma\ell}\right)=M_{0}^{\gamma}\left(\frac{s^{\gamma(L+1)}-1}{s^{\gamma}-1}\right)\simeq M_{0}^{\gamma}s^{\gamma L}\simeq s^{-\frac{\gamma}{\alpha}\log_{s}\epsilon}=\epsilon^{-\gamma/\alpha}.$ (21)
@@ 3.1 Sampling with a single-spatial level surrogate
If the expectation of $u$ is nonzero, the normalized mean squared error (nMSE), denoted ¯$\overline{\mathcal{E}}_{A}^{2}$, is considered. A similar normalization and notation is used for the discretization, interpolation and statistical errors. When the normalizing factor $\left\|\mathbb{E}[u]\right\|_{Z}^{2}$ is not available or computable we approximate it by $\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}$. For instance the nMSE is approximated as
¯$\overline{\mathcal{E}}_{A}^{2}\approx\frac{\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|^{2}_{Z}}{\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}+\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{N\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}\leq\left(\overline{\mathcal{E}}_{\text{Dis}}+\overline{\mathcal{E}}_{\text{Interp}}\right)^{2}+\overline{\mathcal{E}}_{\text{Stat}}^{2}.$
@@
¯$\overline{\mathcal{E}}_{\text{Dis}}\leq\theta_{1}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Interp}}\leq\theta_{2}\epsilon,\qquad\text{ and }\qquad\overline{\mathcal{E}}_{\text{Stat}}=\sqrt{\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{N\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}}\leq\sqrt{\theta}\epsilon,$ (24)
@@
where the subscript “SL-SL”’ represents the use of single-level stochastic collocation and single-level spatial discretization surrogate for online Monte Carlo sampling.
@@ 3.2 Sampling with surrogates with multilevel spatial discretization
and satisfies ${\displaystyle\mathbb{E}\left[A[\widehat{u}_{h}]\right]=\mathbb{E}[\widehat{u}_{L}]}$ and ${\displaystyle\mathbb{V}\left[A[\widehat{u}_{h}]\right]=\sum_{\ell=0}^{L}\frac{\mathbb{V}\left[\widehat{Y}_{\ell}^{(i)}\right]}{N_{\ell}}}$. We remark that for the latter property to hold, the samples $\left(\widehat{u}_{\ell}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right)$ comprising the term $\widehat{Y}_{\ell}^{(i)}$ for each level $\ell$ are drawn independently of each other (i.e. the samples at any given level are not reused for the subsequent levels). However, within each term $\widehat{Y}_{\ell}^{(i)}$, in view of the identity
@@
¯$\overline{\mathcal{E}}_{\text{Dis}}\leq\frac{C_{m}M_{L}^{-\alpha}}{\left\|\mathbb{E}[u_{L}]\right\|_{Z}}\leq\theta_{1}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Interp}}\leq\frac{C_{p}P^{-\nu}}{\left\|\mathbb{E}[u_{L}]\right\|_{Z}}\leq\theta_{2}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Stat}}^{2}=\sum_{\ell=0}^{L}\frac{\mathbb{V}\left[\widehat{Y}_{\ell}\right]}{N_{\ell}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\leq\theta\epsilon^{2},$ (30)
@@
where the superscript “on” stands for the online work, and the subscript “SL-ML” denotes the use of a single level in the parameter space and multiple levels in the discretization space. The goal now is to determine the number $N_{\ell}$ of samples required at every discretization level that will minimize the work requirements while simultaneously yielding a Monte Carlo estimation that satisfies the accuracy threshold ¯$\overline{\mathcal{E}}_{\text{Stat}}\leq\sqrt{\theta}\epsilon$. As shown in [14, 17], this constrained optimization problem yields an estimate for the sample size $N_{\ell}$ given by
@@
Output: Fine-grid level $L$, vector of sample sizes $\boldsymbol{N}=(N_{0}\ldots,N_{L})$ for each discretization level, vector of variance estimations $\boldsymbol{V}=(\mathbb{V}[\widehat{Y}_{0}],\ldots,\mathbb{V}[\widehat{Y}_{L}])$, expectation estimate $A$.
3
4Initialize level $L\leftarrow 0$.
5Initialize vector of sample number corrections $\boldsymbol{dN}\leftarrow\boldsymbol{N}_{\text{old}}$
6while $\sum_{\ell}dN_{\ell}>0$ do
@@
24                  Increase length of the vector $\boldsymbol{N}_{\text{old}}$ by appending 00.
@@
$\sum_{k=0}^{L}s^{\eta k}\simeq\left\{\begin{array}[]{ll}\frac{1}{1-s^{\eta}},&\eta<0,\\ |\log\epsilon|,&\eta=0,\\ \epsilon^{-\eta/\alpha},&\eta>0.\end{array}\right.$ (41)

### v3

# arXiv:2501.08482v3[physics.comp-ph] 01 Mar 2026

@@ 1.1 The Grad-Shafranov free boundary problem with uncertainty
In a magnetic confinement fusion reactor, strong magnetic fields produced by external coils are used to contain a hot plasma in the interior of a vacuum vessel. The confined plasma—a mixture of ionized light atomic nuclei—is heated to induce the merging of the atomic nuclei and the release of a significant amount of energy in a process known as a thermonuclear fusion reaction. During confinement, charged particles in the plasma undergo forces due to the hydrostatic pressure, $p$, and to the magnetic field, $\boldsymbol{B}$, generated by the currents, $\boldsymbol{J}$, flowing through the external coils and through the plasma. A state of equilibrium is reached when the magnetic pressure equals the hydrostatic pressure and is expressed mathematically by requiring that the condition ${\boldsymbol{J}}\times{\boldsymbol{B}}=\nabla p$ holds in the entire space.
If the reactor possesses axial symmetry, the equilibrium condition can then be stated in terms of a nonlinear partial differential equation for a scalar variable known as the poloidal flux, $\psi$. In this study, we will focus on a particular family of axially symmetric reactors known as tokamaks. The cross section of the typical configuration is depicted schematically in Figure 1. A vacuum vessel (solid curved line) contains the plasma (violet region), while an array of coils (solid gray rectangles) generate the confining magnetic field.
@@
$\nabla\,\cdot\,\left(\frac{1}{\mu r}\nabla\psi\right)=J_{\varphi}(\psi,r),$
where $\nabla$ and $\nabla\cdot$ denote the Cartesian gradient and divergence operators in two dimensions respectively; the magnetic permeability $\mu$ is either a function of the magnetic field $\mu=\mu(|\nabla\psi|^{2}/r^{2})$ within any ferromagnetic structure, or equal to the constant permeability of vacuum, $\mu_{0}$, everywhere else. The source $J_{\varphi}$ accounts for the currents present in the system and is known as toroidal current density. This term is defined piecewise and different from zero only in the two following cases:
@@
    Within the plasma (confined within the region $\Omega_{p}$), the current can be represented in terms of the hydrostatic pressure and an additional scalar variable $g$ as
    $J_{\varphi}(\psi,r)=r\frac{d}{d\psi}p(\psi)+\frac{1}{2\,\mu r}\frac{d}{d\psi}g^{2}(\psi).$
Both the pressure and the scalar function $g$ are considered constant along the streamlines of the magnetic field and thus take the forms $p=p(\psi)$ and $g=g(\psi)$. In this study, following the model proposed in [30], we adopt the following forms
$\frac{d}{d\psi}p(\psi)=j_{0}\frac{\beta}{r_{0}}\left(1-\psi_{N}^{a_{1}}\right)^{a_{2}},\qquad\text{ and }\qquad\frac{1}{2}\frac{d}{d\psi}g^{2}(\psi)=j_{0}\mu_{0}r_{0}(1-\beta)\left(1-\psi_{N}^{a_{1}}\right)^{a_{2}},$ (1)
where $\psi_{N}\in[0,1]$ represents a normalization of $\psi$. The parameter $r_{0}$ corresponds to the outer radius of the vacuum chamber, while $a_{1}$ and $a_{2}$ govern the sharpness of the current peaks near the magnetic axis; $\beta$ referred to as poloidal beta, serves as a parameter measuring the ratio between the hydrostatic pressure in the plasma due to temperature (usually called plasma pressure) and the pressure attributed to the external magnetic field (or magnetic pressure), $\mu_{0}$ is the magnetic permeability of vacuum, and $j_{0}$ is introduced as a normalization factor. The region $\Omega_{p}$ circumscribed by the last closed level set of $\psi$ that does not intersect any component of the reactor determines the boundary of the plasma, $\partial\Omega_{p}$, and thus the confinement region. When $\partial\Omega_{p}$ goes through a saddle point of $\psi$, it is referred to as a separatrix, and the saddle point is known as an x-point. Since $\psi$ and $\Omega_{p}$ are not known a priori this leads to a free boundary problem. Moreover, the dependence of the source terms $p(\psi)$ and $g(\psi)$—and possibly even of the magnetic permeability $\mu$—on the solution add further complexity to the Grad-Shafranov free boundary problem.
In practice, all the parameters of the model are susceptible to uncertainties (arising from imperfect measurements, operational variations, engineering tolerances, etc.) that introduce stochasticity into the solution $\psi$ and its derived quantities. In this study we will consider that the uncertainties may affect the parameters appearing in the definition of the source term $J_{\varphi}$; namely, the current intensities of the external coils, and the parameters appearing in (1). To model these uncertainties, we introduce a $d$-dimensional random variable $\boldsymbol{\omega}:=(\omega_{1},\ldots,\omega_{d})$, whose components are independent and uncorrelated, and whose $k$th component $\omega_{k}$ is associated with the uncertainty in the $k$th parameter. We will denote the joint probability density function by $\pi(\boldsymbol{\omega})$ and the parameter space by $\boldsymbol{W}$.
@@
[…]
@@
$\mathbb{E}\left[\psi(\cdot,\boldsymbol{\omega})\right]=\int_{\boldsymbol{W}}\psi(\cdot,\boldsymbol{\omega})\pi(\boldsymbol{\omega})d\boldsymbol{\omega}\,.$ (3)
@@ 1.2 Overview of the article
While our motivating example involves the poloidal flux $\psi$, in order to make the subsequent analysis applicable to a broader class of problems, we introduce a generic state variable $u$ to represent the quantity of interest. In its simplest form, a Monte Carlo estimate of the expectation $\mathbb{E}[u]$ gathers a large number, $N$, of the realizations of $u(\cdot,\boldsymbol{\omega})$ and estimates the expectation by the mean of the sample as
@@ 2.1 Sparse grid stochastic collocation
We now present a brief overview of the sparse grid stochastic collocation method [3, 27, 31, 36] for approximating the solution to (2) with stochastic parameters that will be denoted by $x$ and, for simplicity, will be assumed to be restricted to the $d$-dimensional unit cube $0,1]^{d}$. Assume that for every fixed realization of the parameter $\boldsymbol{\omega}\in\boldsymbol{W}$, $u(\cdot,\boldsymbol{\omega})$ is an element of a function space $Z$. We wish to build an interpolatory approximation of $u(\cdot,\boldsymbol{\omega})$ based on a set of observations at particular values of $\boldsymbol{\omega}$. The stochastic collocation method uses particular values of $\boldsymbol{\omega}$ that give rise to a nested family of grids in the parameter space, where each grid consists of $m_{i}$ values of $\boldsymbol{\omega}$ that we shall call nodes.
To mitigate the so-called curse of dimensionality, the sequence of grids can be built by choosing the nodes according to Smolyak’s rule, which curtails the growth in the number of nodes involved—although at the expense of some accuracy—and leads to a sequence of grids known as sparse grids. The construction process can be summarized as follows. Let
@@ 2.2 Notational considerations
Following the analysis from [21], if $\Omega$ is a bounded Lipschitz domain fully enclosing the component structures of the reactor and all the external coils $\Omega_{c_{i}}$, we will define the solution space to the free boundary problem (2) as
$Z:=\overline{\left\{f:\Omega\rightarrow\mathbb{R}\,\Bigg|\,\int_{\Omega}f^{2}rdrdz<\infty;\,\int_{\Omega}\frac{|\nabla f|^{2}}{r}drdz<\infty;\,f(0,z)=0\right\}\cap C^{0}(\overline{\Omega})}\,.$
Above, the closure is taken with respect to the inner product and the energy norm
$\langle f,g\rangle:=\int_{\Omega}\frac{1}{r}\;\nabla f\cdot\nabla g\;drdz,\qquad\qquad\|f\|_{Z}:=\left(\int_{\Omega}\frac{|\nabla f|^{2}}{r}\;drdz\right)^{1/2}.$
@@
$\left\|u-v\right\|_{L^{2}(\boldsymbol{W},Z)}=\left(\int_{\boldsymbol{W}}\left\|u(\cdot,\boldsymbol{\omega})-v(\cdot,\boldsymbol{\omega})\right\|_{Z}^{2}d\mathbb{P}(\boldsymbol{\omega})\right)^{1/2}=\left(\mathbb{E}\left[\left\|u(\cdot,\boldsymbol{\omega})-v(\cdot,\boldsymbol{\omega})\right\|_{Z}^{2}\right]\right)^{1/2}\,$,
@@ 2.3 Collocation with a single level of spatial discretization
$\mathcal{W}_{\text{SL-SL}}^{\text{off}}=PW\simeq\epsilon^{-\frac{1}{\nu}}\cdot\epsilon^{-\frac{\gamma}{\alpha}}=\epsilon^{-\frac{1}{\nu}-\frac{\gamma}{\alpha}},$
@@ 2.4 Collocation with multilevel spatial discretization
$\leq\,$ $\|u-u_{L}\|_{L^{2}(\boldsymbol{W},Z)}+\sum_{\ell=0}^{L}\left\|Y_{\ell}-\widehat{Y}_{\ell}\right\|_{L^{2}(\boldsymbol{W},Z)}$ $\qquad{\text{\small(Telescopic expansion for $u_{L}$ and $\widehat{u}_{L}$ )}},$
@@
where $\left\lceil\cdot\right\rceil$ denotes the ceiling function.
@@
$\sum_{\ell=0}^{L}W_{\ell}\simeq\sum_{\ell=0}^{L}{M_{\ell}^{\gamma}}=M_{0}^{\gamma}\left(\sum_{\ell=0}^{L}s^{\gamma\ell}\right)=M_{0}^{\gamma}\left(\frac{s^{\gamma(L+1)}-1}{s^{\gamma}-1}\right)\simeq M_{0}^{\gamma}s^{\gamma L}\simeq s^{-\frac{\gamma}{\alpha}\log_{s}\epsilon}=\epsilon^{-\gamma/\alpha}.$
@@ 3.1 Sampling with a single-spatial level surrogate
If the expectation of $u$ is nonzero, the normalized mean squared error (nMSE), denoted $\overline{\mathcal{E}}_{A}^{2}$, is considered. A similar normalization and notation is used for the discretization, interpolation and statistical errors. When the normalizing factor $\left\|\mathbb{E}[u]\right\|_{Z}^{2}$ is not available or computable we approximate it by $\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}$. For instance the nMSE is approximated as
$\overline{\mathcal{E}}_{A}^{2}\approx\frac{\left\|\mathbb{E}[u]-\mathbb{E}\left[\widehat{u}_{h}\right]\right\|^{2}_{Z}}{\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}+\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{N\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}\leq\left(\overline{\mathcal{E}}_{\text{Dis}}+\overline{\mathcal{E}}_{\text{Interp}}\right)^{2}+\overline{\mathcal{E}}_{\text{Stat}}^{2}.$
@@
$\overline{\mathcal{E}}_{\text{Dis}}\leq\theta_{1}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Interp}}\leq\theta_{2}\epsilon,\qquad\text{ and }\qquad\overline{\mathcal{E}}_{\text{Stat}}=\sqrt{\frac{\mathbb{V}\left[\widehat{u}_{h}\right]}{N\left\|\mathbb{E}[u_{h}]\right\|_{Z}^{2}}}\leq\sqrt{\theta}\epsilon,$ (22)
@@
where the subscript “SL-SL” represents the use of single-level stochastic collocation and single-level spatial discretization surrogate for online Monte Carlo sampling.
@@ 3.2 Sampling with surrogates with multilevel spatial discretization
and satisfies $\mathbb{E}\left[A[\widehat{u}_{h}]\right]=\mathbb{E}[\widehat{u}_{L}]$} and $\mathbb{V}\left[A[\widehat{u}_{h}]\right]=\sum_{\ell=0}^{L}\frac{\mathbb{V}\left[\widehat{Y}_{\ell}^{(i)}\right]}{N_{\ell}}$}. We remark that for the latter property to hold, the samples $\left(\widehat{u}_{\ell}^{(i)}-\widehat{u}_{\ell-1}^{(i)}\right)$ comprising the term $\widehat{Y}_{\ell}^{(i)}$ for each level $\ell$ are drawn independently of each other (i.e. the samples at any given level are not reused for the subsequent levels). However, within each term $\widehat{Y}_{\ell}^{(i)}$, in view of the identity
@@
$\overline{\mathcal{E}}_{\text{Dis}}\leq\frac{C_{m}M_{L}^{-\alpha}}{\left\|\mathbb{E}[u_{L}]\right\|_{Z}}\leq\theta_{1}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Interp}}\leq\frac{C_{p}P^{-\nu}}{\left\|\mathbb{E}[u_{L}]\right\|_{Z}}\leq\theta_{2}\epsilon,\qquad\overline{\mathcal{E}}_{\text{Stat}}^{2}=\sum_{\ell=0}^{L}\frac{\mathbb{V}\left[\widehat{Y}_{\ell}\right]}{N_{\ell}\left\|\mathbb{E}[u_{L}]\right\|_{Z}^{2}}\leq\theta\epsilon^{2},$ (28)
@@
where the superscript “on” stands for the online work, and the subscript “SL-ML” denotes the use of a single level in the parameter space and multiple levels in the discretization space. The goal now is to determine the number $N_{\ell}$ of samples required at every discretization level that will minimize the work requirements while simultaneously yielding a Monte Carlo estimation that satisfies the accuracy threshold $\overline{\mathcal{E}}_{\text{Stat}}\leq\sqrt{\theta}\epsilon$. As shown in [14, 17], this constrained optimization problem yields an estimate for the sample size $N_{\ell}$ given by
@@
31ex
Output: Fine-grid level $L$, vector of sample sizes $\boldsymbol{N}=(N_{0}\ldots,N_{L})$ for each discretization level, vector of variance estimations $\boldsymbol{V}=(\mathbb{V}[\widehat{Y}_{0}],\ldots,\mathbb{V}[\widehat{Y}_{L}])$, expectation estimate AA.
4
51ex
6
7Initialize level $L\leftarrow 0$.
8Initialize vector of sample number corrections $\boldsymbol{dN}\leftarrow\boldsymbol{N}_{\text{old}}$
9while $\sum_{\ell}dN_{\ell}>0$ do
@@
27        Increase length of the vector $\boldsymbol{N}_{\text{old}}$ by appending 0.
@@
$\sum_{k=0}^{L}s^{\eta k}\simeq\left\{\begin{array}[]{ll}\frac{1}{1-s^{\eta}},&\eta<0,\\ |\log\epsilon|,&\eta=0,\\ \epsilon^{-\eta/\alpha},&\eta>0.\end{array}\right$.